# Beyond ReAct: A Planner-Centric Framework for Complex Tool-Augmented LLM Reasoning

### v1

@@ Introduction
Large Language Models (LLMs) have demonstrated remarkable prowess in language processing, yet their inherent knowledge is static and they lack direct interaction with the external world. A transformative solution is tool augmentation, which empowers LLMs to act as autonomous agents by calling external APIs and tools (Schick et al. 2023; Patil et al. 2024). This paradigm shifts LLMs from being mere text generators to active problem-solvers capable of tackling complex, real-world tasks, a trend exemplified by systems like HuggingGPT (Shen et al. 2023) and the rise of multi-agent frameworks (Hong et al. 2023). The evaluation of these agentic capabilities has itself become a major research thrust, with new benchmarks like AgentBench (Liu et al. 2023) setting the stage for more rigorous assessment.
@@
The dominant approach for orchestrating tool usage relies on reactive, step-by-step reasoning frameworks like ReAct (Yao et al. 2023b), often augmented by self-reflection techniques (Shinn et al. 2023). However, this paradigm suffers from inherent local optimization traps due to its incremental decision-making process. While potentially effective for simple queries, its reactive nature often falters on complex tasks. As illustrated in Fig. 1, complex tasks involve more intricate dependencies than simple ones. To mitigate this, advanced methods employ tree search (Yao et al. 2023a) or explicit planning prompts (Wang et al. 2023) to explore multiple paths. While these methods improve upon purely linear reasoning, they fail to architecturally resolve the fundamental bottleneck of local optimization. Critically, they remain focused on finding an optimal sequential path, overlooking opportunities for parallelism inherent in complex workflows. Moreover, the computational overhead of constructing and traversing search trees, requiring numerous LLM calls, can be prohibitive.
@@
To address this fundamental challenge, we adopt a framework that decouples planning from execution (Gao et al. 2023; Chen et al. 2022). Our approach employs a dedicated Planner model, trained to translate complex user queries into an execution plan structured as a DAG. In this DAG, nodes represent selected tools, and directed edges capture dependency relationships between them. By optimizing the Planner to improve the accuracy of both node selection (tool choices) and edge prediction (dependency relationships) in the DAG, we achieve more effective plans for complex queries, improving workflow execution compared to sequential models.
@@
Realizing such a global planner presents two significant hurdles. First, there is a lack of large-scale, structured data required to train a model to generate complex plans, a challenge addressed by recent large-scale tool-use benchmarks like ToolBench (Qin et al. 2023) and API-Bank (Li et al. 2023b). Second, evaluating the quality of a plan is non-trivial. To overcome these, we introduce ComplexTool-Plan, a new, large-scale benchmark with graded difficulty levels. It serves as both the training ground and the evaluation standard for sophisticated planning agents. We then devise a two-stage hybrid training strategy for our Planner, combining Supervised Fine-Tuning (SFT) with subsequent refinement via Reinforcement Learning (RL) (Ouyang et al. 2022), a strategy that aligns with recent findings on using self-play and RL to elicit stronger reasoning (Chen et al. 2024; Wei et al. 2025).
@@
Our extensive experiments validate the efficacy of our decoupled approach. On ComplexTool-Plan, our Planner substantially outperforms leading baselines. Crucially, when our Planner is paired with a powerful executor model like GPT-4o, the integrated system establishes a new state-of-the-art (SOTA) on the widely recognized end-to-end benchmark, StableToolBench (Guo et al. 2024), underscoring the practical superiority of our framework.
@@
We survey the rapidly evolving landscape of tool-augmented LLMs along three primary axes: the evolution of tool-calling paradigms, RL-based policy optimization, and the development of evaluation benchmarks.
@@
The paradigms for integrating LLMs with external tools have evolved from the interleaved ”Reasoning-Acting” steps of ReAct (Yao et al. 2023b) to more reliable, natively supported structured API calls. A significant branch of research leverages code as a universal tool interface, with works like Program-of-Thought (PoT) (Chen et al. 2022) and PAL (Gao et al. 2023) offloading complex computations to code interpreters. This trend has produced highly specialized agents like SWE-agent (Yang et al. 2024), which can autonomously resolve software engineering tasks. For instance, advanced models are increasingly required for challenging real-world applications like multi-modal fake news detection, which involves reasoning over diverse data from various sources and domains (Tong et al. 2025; Lu, Tong, and Ye 2025).
Recent work is also moving beyond general-purpose models towards those specifically fine-tuned for tool use. For instance, Toolformer (Schick et al. 2023) demonstrates that a language model can teach itself to use tools by learning to insert beneficial API calls into text, thus significantly improving its ability to handle complex, tool-dependent instructions in a zero-shot manner. Concurrently, the interaction paradigm is expanding to visual environments, with powerful GUI agents like CogAgent (Hong et al. 2024) demonstrating the ability to understand and interact with graphical user interfaces. Beyond reactive or single-agent frameworks, Li et al. (Li et al. 2025) introduce an AI Search Paradigm that formalizes reasoning as coordinated planning among multiple LLM-based agents, while Chen et al. (Chen et al. 2025) propose a proactive multi-agent orchestration framework for complex information-seeking tasks. Our work contributes to this evolution by focusing on the upfront, holistic planning of tool interactions, a critical component for efficient orchestration.
@@
While Supervised Fine-Tuning (SFT) can teach models to mimic tool-use trajectories, Reinforcement Learning (RL) is superior for strategic decision-making in unseen scenarios. The latest SOTA models, such as Llama 3 (Dubey et al. 2024), heavily rely on advanced RLHF techniques to refine their policy. Building on foundational work demonstrating that RL can enhance reasoning (Guo et al. 2025), recent research has shown that RL teaches not just the ”syntax” of tool calling, but the ”strategy” (Feng et al. 2025; Jin et al. 2025).
The field is advancing towards more sophisticated, outcome-driven optimization loops. For instance, AlphaCodium (Ridnik, Kredo, and Friedman 2024) employs a test-based, multi-stage iterative process. Similarly, influential frameworks like Tree of Thoughts (ToT) (Yao et al. 2023a), which explores multiple reasoning paths, and Reflexion (Shinn et al. 2023), which leverages self-reflection, highlight the trend towards more deliberate and verifiable reasoning. Our hierarchical reward function aligns with this trend by providing a structured, outcome-driven signal to teach the model a robust planning strategy.
@@
The evaluation of LLM tool-calling capabilities relies on increasingly sophisticated benchmarks, which have evolved from early interactive environments like ALFWorld (Shridhar et al. 2020) and WebShop (Yao et al. 2022). In recent years, benchmarks have grown in scale and realism, with examples like ToolBench (Qin et al. 2023) and its more stable successor, StableToolBench (Guo et al. 2024). The focus has also expanded towards greater realism, as seen in the web environment of WebArena (Zhou et al. 2023), and more challenging tasks, as in the GAIA benchmark (Mialon et al. 2023).
To support this research, new platforms are emerging that focus on specific aspects like evaluating the correctness of API calls in API-Bank (Li et al. 2023b) and facilitating multi-agent collaboration in open-source frameworks like MetaGPT (Hong et al. 2023). Our work contributes a new benchmark, ComplexTool-Plan, which fills a specific gap in this landscape by targeting the evaluation of an agent’s ability to generate complex, non-linear plans.
@@
To ensure a stable and efficient RL phase, we curate the training data to focus on problems at the frontier of the model’s current capabilities (Chen et al. 2024). We use the SFT-trained model as a filter: tasks that the model already consistently solves (offering no learning signal) or consistently fails (intractably hard) are excluded from the RL training set. This curation of high-variance instances, where the outcome is uncertain, concentrates the training on the most informative samples, preventing policy degradation and fostering robust optimization.
@@
To move beyond SFT’s limitations, we use Reinforcement Learning (RL). This, however, requires a nuanced reward signal to properly evaluate complex DAGs, distinguishing structural from strategic errors. We address this by introducing a Hierarchical Reward Function $R(y)$. To optimize this complex, multi-dimensional reward, we utilize the Group Relative Policy Optimization (GRPO) algorithm (Shao et al. 2024), which is specifically designed for such scenarios and ensures robust policy improvement.
@@
We evaluate our Planner on ComplexTool-Plan, our benchmark constructed from a comprehensive library of 4,535 tool APIs sourced from ModelScope (Li et al. 2023a). The benchmark features three difficulty levels (Easy, Medium, and Hard) and provides a curated SFT dataset of 3,000 instances. For RL training, the dataset is filtered to comprise 787 instances, which are subsequently divided into training and test sets using an 8:2 ratio. Further details are available in Fig. 3.
@@
To evaluate the practical effectiveness of our framework, we use StableToolBench (Guo et al. 2024) for end-to-end evaluation. StableToolBench is an enhanced version of the test set from ToolBench (Qin et al. 2023), specifically designed to mitigate execution instability via a caching system and API simulators. It categorizes test cases across two axes: tool generalization and scenario complexity.
@@
  • Open-Source Frameworks: For end-to-end evaluation on StableToolBench (Guo et al. 2024), we compare against established methods applied to the ToolLLaMA model (Qin et al. 2023). This includes the reactive framework ReAct, the planning-based method DFSDT, and a highly relevant concurrent work, LLMCompiler (Kim et al. 2024).
@@
While iterative paradigms like DTA-Llama (Zhu et al. 2025) report a higher SoPR, their success is largely driven by a multi-turn framework that allows for in-process error correction and adaptation. In contrast, our non-iterative, plan-then-execute paradigm has only one chance to generate a correct global plan upfront. This architectural choice prioritizes predictability and planning quality over the adaptive resilience of iterative feedback loops.

### v2

@@ Introduction
Large Language Models (LLMs) have demonstrated remarkable prowess in language processing, yet their inherent knowledge is static and they lack direct interaction with the external world. A transformative solution is tool augmentation, which empowers LLMs to act as autonomous agents by calling external APIs and tools (schick2023toolformer; patil2024gorilla). This paradigm shifts LLMs from being mere text generators to active problem-solvers capable of tackling complex, real-world tasks, a trend exemplified by systems like HuggingGPT (shen2023hugginggpt) and the rise of multi-agent frameworks (hong2023metagpt). The evaluation of these agentic capabilities has itself become a major research thrust, with new benchmarks like AgentBench (liu2023agentbench) setting the stage for more rigorous assessment.
@@
The dominant approach for orchestrating tool usage relies on reactive, step-by-step reasoning frameworks like ReAct (yao2023react), often augmented by self-reflection techniques (shinn2023reflexion). However, this paradigm suffers from inherent local optimization traps due to its incremental decision-making process. While potentially effective for simple queries, its reactive nature often falters on complex tasks. As illustrated in Fig. 1, complex tasks involve more intricate dependencies than simple ones. To mitigate this, advanced methods employ tree search (yao2023tree) or explicit planning prompts (wang2023plan) to explore multiple paths. While these methods improve upon purely linear reasoning, they fail to architecturally resolve the fundamental bottleneck of local optimization. Critically, they remain focused on finding an optimal sequential path, overlooking opportunities for parallelism inherent in complex workflows. Moreover, the computational overhead of constructing and traversing search trees, requiring numerous LLM calls, can be prohibitive.
@@
To address this fundamental challenge, we adopt a framework that decouples planning from execution (gao2023pal; chen2022program). Our approach employs a dedicated Planner model, trained to translate complex user queries into an execution plan structured as a DAG. In this DAG, nodes represent selected tools, and directed edges capture dependency relationships between them. By optimizing the Planner to improve the accuracy of both node selection (tool choices) and edge prediction (dependency relationships) in the DAG, we achieve more effective plans for complex queries, improving workflow execution compared to sequential models.
@@
Realizing such a global planner presents two significant hurdles. First, there is a lack of large-scale, structured data required to train a model to generate complex plans, a challenge addressed by recent large-scale tool-use benchmarks like ToolBench (qin2023toolllm) and API-Bank (li2023api). Second, evaluating the quality of a plan is non-trivial. To overcome these, we introduce ComplexTool-Plan, a new, large-scale benchmark with graded difficulty levels. It serves as both the training ground and the evaluation standard for sophisticated planning agents. We then devise a two-stage hybrid training strategy for our Planner, combining Supervised Fine-Tuning (SFT) with subsequent refinement via Reinforcement Learning (RL) (ouyang2022training), a strategy that aligns with recent findings on using self-play and RL to elicit stronger reasoning (chen2024self; wei2025igniting).
@@
Our extensive experiments validate the efficacy of our decoupled approach. On ComplexTool-Plan, our Planner substantially outperforms leading baselines. Crucially, when our Planner is paired with a powerful executor model like GPT-4o, the integrated system establishes a new state-of-the-art (SOTA) on the widely recognized end-to-end benchmark, StableToolBench (guo2024stabletoolbench), underscoring the practical superiority of our framework.
@@
We survey the rapidly evolving landscape of tool-augmented LLMs(li2025rankexpert; li2025rankelectra) along three primary axes: the evolution of tool-calling paradigms, RL-based policy optimization, and the development of evaluation benchmarks.
@@
The paradigms for integrating LLMs with external tools have evolved from the interleaved ”Reasoning-Acting” steps of ReAct (yao2023react) to more reliable, natively supported structured API calls. A significant branch of research leverages code as a universal tool interface, with works like Program-of-Thought (PoT) (chen2022program) and PAL (gao2023pal) offloading complex computations to code interpreters. This trend has produced highly specialized agents like SWE-agent (yang2024swe), which can autonomously resolve software engineering tasks. For instance, advanced models are increasingly required for challenging real-world applications like multi-modal fake news detection, which involves reasoning over diverse data from various sources and domains (tong2025dapt; lu2025dammfnd).
Recent work is also moving beyond general-purpose models towards those specifically fine-tuned for tool use(zhao2025turatoolaugmentedunifiedretrieval). For instance, Toolformer (schick2023toolformer) demonstrates that a language model can teach itself to use tools by learning to insert beneficial API calls into text, thus significantly improving its ability to handle complex, tool-dependent instructions in a zero-shot manner. Concurrently, the interaction paradigm is expanding to visual environments, with powerful GUI agents like CogAgent (hong2024cogagent) demonstrating the ability to understand and interact with graphical user interfaces. Beyond reactive or single-agent frameworks, Li et al. (li2025towards) introduce an AI Search Paradigm that formalizes reasoning as coordinated planning among multiple LLM-based agents, while Chen et al. (chen2025multi) propose a proactive multi-agent orchestration framework for complex information-seeking tasks. Our work contributes to this evolution by focusing on the upfront, holistic planning of tool interactions, a critical component for efficient orchestration.
@@
While Supervised Fine-Tuning (SFT) can teach models to mimic tool-use trajectories, Reinforcement Learning (RL) is superior for strategic decision-making in unseen scenarios. The latest SOTA models, such as Llama 3 (dubey2024llama), heavily rely on advanced RLHF techniques to refine their policy. Building on foundational work demonstrating that RL can enhance reasoning (guo2025deepseek), recent research has shown that RL teaches not just the ”syntax” of tool calling, but the ”strategy” (feng2025retool; jin2025search).
The field is advancing towards more sophisticated, outcome-driven optimization loops. For instance, AlphaCodium (ridnik2024code) employs a test-based, multi-stage iterative process. Similarly, influential frameworks like Tree of Thoughts (ToT) (yao2023tree), which explores multiple reasoning paths, and Reflexion (shinn2023reflexion), which leverages self-reflection, highlight the trend towards more deliberate and verifiable reasoning. Our hierarchical reward function aligns with this trend by providing a structured, outcome-driven signal to teach the model a robust planning strategy.
@@
The evaluation of LLM tool-calling capabilities relies on increasingly sophisticated benchmarks, which have evolved from early interactive environments like ALFWorld (shridhar2020alfworld) and WebShop (yao2022webshop). In recent years, benchmarks have grown in scale and realism, with examples like ToolBench (qin2023toolllm) and its more stable successor, StableToolBench (guo2024stabletoolbench). The focus has also expanded towards greater realism, as seen in the web environment of WebArena (zhou2023webarena), and more challenging tasks, as in the GAIA benchmark (mialon2023gaia).
To support this research, new platforms are emerging that focus on specific aspects like evaluating the correctness of API calls in API-Bank (li2023api) and facilitating multi-agent collaboration in open-source frameworks like MetaGPT (hong2023metagpt). Our work contributes a new benchmark, ComplexTool-Plan, which fills a specific gap in this landscape by targeting the evaluation of an agent’s ability to generate complex, non-linear plans.
@@
To ensure a stable and efficient RL phase, we curate the training data to focus on problems at the frontier of the model’s current capabilities (chen2024self). We use the SFT-trained model as a filter: tasks that the model already consistently solves (offering no learning signal) or consistently fails (intractably hard) are excluded from the RL training set. This curation of high-variance instances, where the outcome is uncertain, concentrates the training on the most informative samples, preventing policy degradation and fostering robust optimization.
@@
To move beyond SFT’s limitations, we use Reinforcement Learning (RL). This, however, requires a nuanced reward signal to properly evaluate complex DAGs, distinguishing structural from strategic errors. We address this by introducing a Hierarchical Reward Function $R(y)$. To optimize this complex, multi-dimensional reward, we utilize the Group Relative Policy Optimization (GRPO) algorithm (shao2024deepseekmath), which is specifically designed for such scenarios and ensures robust policy improvement.
@@
We evaluate our Planner on ComplexTool-Plan, our benchmark constructed from a comprehensive library of 4,535 tool APIs sourced from ModelScope (li2023modelscope). The benchmark features three difficulty levels (Easy, Medium, and Hard) and provides a curated SFT dataset of 3,000 instances. For RL training, the dataset is filtered to comprise 787 instances, which are subsequently divided into training and test sets using an 8:2 ratio. Further details are available in Fig. 3.
@@
To evaluate the practical effectiveness of our framework, we use StableToolBench (guo2024stabletoolbench) for end-to-end evaluation. StableToolBench is an enhanced version of the test set from ToolBench (qin2023toolllm), specifically designed to mitigate execution instability via a caching system and API simulators. It categorizes test cases across two axes: tool generalization and scenario complexity.
@@
  • Open-Source Frameworks: For end-to-end evaluation on StableToolBench (guo2024stabletoolbench), we compare against established methods applied to the ToolLLaMA model (qin2023toolllm). This includes the reactive framework ReAct, the planning-based method DFSDT, and a highly relevant concurrent work, LLMCompiler (kim2024llm).
@@
While iterative paradigms like DTA-Llama (zhu2025divide) report a higher SoPR, their success is largely driven by a multi-turn framework that allows for in-process error correction and adaptation. In contrast, our non-iterative, plan-then-execute paradigm has only one chance to generate a correct global plan upfront. This architectural choice prioritizes predictability and planning quality over the adaptive resilience of iterative feedback loops.